# kNNSampler: Stochastic Imputations for Recovering Missing Value Distributions

### v1

@@ Abstract
We study a missing-value imputation method, termed kNNSampler, that imputes a given unit’s missing response by randomly sampling from the observed responses of the $k$ most similar units to the given unit in terms of the observed covariates. This method can sample unknown missing values from their distributions, quantify the uncertainties of missing values, and be readily used for multiple imputation. Unlike popular kNNImputer, which estimates the conditional mean of a missing response given an observed covariate, kNNSampler is theoretically shown to estimate the conditional distribution of a missing response given an observed covariate. Experiments demonstrate its effectiveness in recovering the distribution of missing values. The code for kNNSampler is made publicly available. [^1] https://github.com/SAP/knn-sampler
@@ 1 Introduction
Missing values occur in real-world datasets for various reasons, such as non-response in surveys and sensor failures. Imputation — filling in missing values with their estimates — is a common preprocessing step used to address missing data. Over the decades, various imputation methods have been proposed, ranging from simple statistical techniques to machine learning algorithms (e.g., Rubin,, 1976; Schafer,, 1997; Schafer and Graham,, 2002; Little and Rubin,, 2002; Mattei and Frellsen,, 2019; Enders,, 2022).
kNNImputer (Troyanskaya et al.,, 2001) is one of the most widely used imputation methods, owing to its simplicity and availability in popular software packages such as scikit-learn [^2] https://scikit-learn.org/stable/modules/generated/sklearn.impute.KNNImputer.html (Pedregosa et al.,, 2011). It imputes a missing response variable (e.g., customer satisfaction level) of a given unit (e.g., a customer) as the average of the observed responses of the $k$ most similar units to the given unit in terms of observed covariates (e.g., age, gender, occupation). This is to predict the missing response by $k$ nearest neighbours (kNN) regression (Stone,, 1977) so the imputation is an estimate of the conditional expectation of the missing response given a covariate. The method has been widely used in science and engineering, and many extensions have been proposed (e.g., Garc´ıa-Laencina et al.,, 2009; Tutz and Ramzan,, 2015; De Silva and Perera,, 2016; Huang et al.,, 2017; Faisal and Tutz,, 2021).
@@
An issue of kNNImputer, shared by other regression-based imputers, is that the distribution of imputations can be significantly different from the distribution of true (hidden) missing values. This is because, as mentioned, an imputation of kNNImputer is an estimate of the conditional expectation of a missing response, thus tending to be a deterministic function of the covariate. As a result, the distribution of imputed responses is concentrated around the regression curve, even when the distribution of missing responses has large variability. This is illustrated in Figures 1 and 2, where the true conditional distribution of a missing response is bimodal when the covariate is small, but the distribution of imputations is unimodal and many imputations take values never realized by the true missing values. A substantial bias can occur in an analysis of such a distorted imputed dataset, for example, when estimating the variance, quantiles and modes in the population.
@@
kNNSampler can be interpreted as an instance of hot deck, classic imputation methods widely used in practice for socio-economic and public health surveys, including the U.S. Census Bureau’s Current Population Survey and the National Center for Education Statistics (Andridge and Little,, 2010, e.g.,). In a hot deck method, a missing value of a given unit is imputed as one of the response values of the units belonging to the same “adjustment cell” as the given unit. The method is called random hot deck if the imputation is selected randomly from the adjustment cell; it is called nearest-neighbour hot deck if nearest neighbours define the adjustment cell (Little and Rubin,, 2002, Example 4.9). kNNSampler is thus essentially a nearest-neighbour random hot deck method. However, while classic and widely used, hot deck methods have not been well established theoretically (Andridge and Little,, 2010).
@@
  • We derive an error bound between the kNN and true conditional distributions for any given, fixed covariate, in terms of the number $n$ of observed response-covariate pairs, the number $k$ of the nearest neighbours, and other problem-specific constants. The error is measured by the maximum mean discrepancy (MMD) (Gretton et al.,, 2012), a distance metric on probability distributions that metrizes the weak convergence (Simon-Gabriel et al.,, 2023), between the kNN and true conditional distributions. It holds under a Lipschitz condition that the response’s conditional distribution changes smoothly when the covariate changes continuously. A consequence of the bound is the statistical consistency of the kNN conditional distribution, in that the error decreases to zero as the sample size $n$ goes to infinity, if the number $k$ of nearest neighbours increases to infinity at a rate slower than $n$. This offers a theoretical foundation of the kNNSampler and thus the nearest-neighbour random hot deck.
  • To derive the bound, we analyze the mean embedding of the kNN conditional distribution in a reproducing kernel Hilbert space (RKHS) as a novel estimator of the mean embedding of the true conditional distribution, known as conditional mean embedding (Muandet et al.,, 2017, Chapter 4), which is the RKHS-valued regression function (Grünewälder et al.,, 2012). The RKHS distance between these two embeddings is equivalent to the MMD between the kNN and true conditional distributions. Our bound leads to the consistency and convergence rates for the novel kNN-based estimator of the conditional mean embedding.
  • Our analysis extends the error analysis by Kpotufe, (2011) on real-valued kNN regression to RKHS-valued regression in which the response variable is infinite-dimensional. As a byproduct, we prove that the required sample size to attain a given level of precision increases exponentially not with the covariate’s ambient dimension but with the intrinsic dimension of the covariate distribution. Therefore, the kNNSampler may not be severely affected by the curse of dimensionality if the covariate distribution has a low intrinsic dimension.
This paper is organised as follows. We describe the proposed approach in Section 2, its theory in Section 3, and experiments in Section 4.
@@ 2.1 Setting
This assumption implies that the probability of a unit missing its response is determined by the unit’s covariate and is not affected by the response. Therefore, it is an instance of the Missing-At-Random (MAR) assumption (Rubin,, 1976). In the special case where the covariate distributions for the two cases are the same, Q(x~)=P(x~)Q($\tilde${x})=P($\tilde${x}), the assumption can be interpreted as the Missing-Completely-At-Random (MCAR) assumption where missingness occurs completely randomly.
Under this setup, missing responses may be imputed by estimating the unknown conditional distribution $P(y|x)$ of a response given a covariate, and sampling from the estimated conditional distribution. This is what the kNNSampler does.
@@ 2.2 Issue with kNNImputer and Regression-based Imputers
Before describing the proposed kNNSampler, we discuss an issue with the widely used kNNImputer (Troyanskaya et al.,, 2001) and other regression-based imputation methods.
@@
kNNImputer (Troyanskaya et al.,, 2001) imputes the missing response $\tilde{y}_{i,{\rm miss}}$ of the unit with observed covariate $\tilde{x}_{i}$ as the average of the observed responses $y_{j_{1}},\dots,y_{j_{k}}$ of its $k$-nearest neighbors $x_{j_{1}},\dots,x_{j_{k}}$:
@@
This is kNN regression (e.g., Györfi et al.,, 2002) and thus estimates the conditional mean of the missing response $\tilde{y}_{i,{\rm miss}}$ given the observed covariate $\tilde{x}_{i}$:
@@
The number of nearest neighbors $k$ is a hyperparameter of kNNSampler. The theoretical and empirical results below indicate that $k$ should not be fixed to a prespecified value (e.g., $k=5$), and should be chosen depending on the available data. One way is to perform cross-validation for kNN regression on the data $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ and select $k$ among candidate values that minimizes the mean-square error on held-out observed responses, averaged over different training-validation splits. In particular, the present work uses Leave-One-Out Cross-Validation (LOOCV) using the fast computation method recently proposed by Kanagawa, (2024).
@@ 2.5 Multiple Imputation with kNNSampler
An analysis can then be made based on the standard procedure of multiple imputation (Rubin,, 1987).
@@ 3 Theory
We describe a theory for kNNSampler’s conditional distribution (6) as an estimator of the true conditional distribution. We shall show that, as the number $k$ of nearest neighbors increases at an approximate rate as the increase of the number $n$ of observed covariate-response pairs, the kNN conditional distribution converges to the true one in the Maximum Mean Discrepancy (MMD) (Gretton et al.,, 2012), which implies the convergence in distribution (Sriperumbudur et al.,, 2010, Section 5). We prove this by adapting the proof of the convergence rates of real-valued kNN regression by Kpotufe, (2011, Theorem 1) to Hilbert space-valued kNN regression. [^4] Hilbert space-valued kNN regression was also analyzed in Lian, (2011), but their results are not directly applicable to our case. This is because Lian, (2011) assumes that Hilbert space-valued noises are independent of input variables, but this assumption is too strong in our case.
We use the framework of kernel mean embedding (Muandet et al.,, 2017) in which every probability distribution is represented as a distinct point in an infinite-dimensional feature space known as a reproducing kernel Hilbert space (RKHS). The true and kNN conditional distributions are represented as points in an RKHS, and the distance between them, which is the MMD, quantifies the estimation error. An upper bound on this distance is obtained in terms of the sample size, the number of nearest neighbours, and other relevant quantities.
@@ 3.1 RKHS Embeddings of Conditional Distributions
This is called reproducing kernel of the RKHS. The RKHS and the reproducing kernel are one-to-one, so an RKHS can be induced by defining a kernel. For example, if $\mathcal{Y}=\mathbb{R}^{p}$, the Gaussian kernel $\ell(y,y^{\prime})=\exp(-\alpha\|y-y^{\prime}\|^{2})$ for $\alpha>0$ is the reproducing kernel of a certain RKHS $\mathcal{H}$, and there exists an infinite-dimensional feature map $\Phi$ that induces the Gaussian kernel as (7). See e.g. Steinwart and Christmann, (2008); Kanagawa et al., (2025) for details on RKHSs.
@@
In this case, the RKHS is called characteristic (Sriperumbudur et al.,, 2010). For example, Gaussian, Matérn and Laplace kernels induce characteristics RKHSs.
@@
The mean embedding of the conditional distribution is known as conditional mean embedding (Song et al.,, 2009; 2013) and its estimator based on a regularized least-squares algorithm has been studied extensively (e.g., Grünewälder et al.,, 2012; Li et al.,, 2022; 2024). The mean embedding of the kNN conditional distribution in (8) is a new estimator of the conditional mean embedding. Its analysis below is thus a new contribution to the RKHS literature and may be of independent interest.
@@ 3.2 Assumptions
We describe key assumptions for the analysis, which follow Kpotufe, (2011) with appropriate modifications.
The conditional mean embedding in (8) is the conditional expectation of the response feature vector $\Phi(y)$ given a covariate $x\in\mathcal{X}$; thus, it is the RKHS-valued regression function (Grünewälder et al.,, 2012). We assume that the map from a covariate $x$ to the conditional mean embedding $\Phi(P(\cdot\mid x))$ is smooth in the sense that it is Lipschitz continuous.
@@
This assumption is satisfied, for example, if $\mathcal{X}=\mathbb{R}^{p}$ with $p\geq 1$, in which case $\mathcal{V}_{\mathcal{B}}\leq p+2$ (e.g., Mohri et al.,, 2018, Exercise 3.17).
@@ 3.3 Error Bounds and Convergence Rates
Under the above assumptions, the distance between the true and kNN conditional distributions can be upper-bounded as follows. The proof, provided in Appendix A, is an adaptation of the proof of Kpotufe, (2011, Theorem 1), which is an upper error bound on real-valued kNN regression, to our setting of RKHS-valued regression.
@@
$\left\|\Phi(P(\cdot\mid x))-\Phi(\hat{P}(\cdot\mid x))\right\|_{\mathcal{H}}^{2}\leq C_{1}\frac{\ln(n)}{k}+C_{2}\left(\frac{k}{n}\right)^{d/2},$ (12)
@@
On the other hand, if $k$ is fixed to a constant value (e.g., $k=1$), the variance term does not decrease even if the sample size increases. These observations are well known for real-valued kNN regression (e.g., Györfi et al.,, 2002).
@@
The above consistency (13) implies the convergence in distribution (or weak convergence) of the kNN conditional distribution $\hat{P}(\cdot\mid x)$ to the true one $P(\cdot\mid x)$ if the response space $\mathcal{Y}$ is a compact metric space (e.g., $\mathcal{Y}$ is a bounded closed subset of an Euclidean space) and $\mathcal{H}$ is a universal RKHS [^6] An RKHS $\mathcal{H}$ consisting of functions on a metric set $\mathcal{Y}$ is called universal if any continuous bounded function $f:\mathcal{Y}\to\mathbb{R}$ can be approximated arbitrarily well in terms of the supremum norm by functions in $\mathcal{H}$., such as the RKHSs of Gaussian, Matérn and Laplace kernels (Sriperumbudur et al.,, 2010, Theorem 23); see Simon-Gabriel et al., (2023) for more generic conditions. That is, under these conditions, the expectation of any continuous bounded function $f:\mathcal{Y}\to\mathbb{R}$ under the kNN distribution $\hat{P}(\cdot\mid x)$ converges to the expectation under the true distribution $P(\cdot\mid x)$:
@@
The rate (14) shows that the required sample size $n$ to attain a desired error level increases exponentially with respect to the intrinsic dimension $d$ of the covariate distribution $P(x)$, not the ambient dimension of the input space $\mathcal{X}$, which is captured by the VC dimension $\mathcal{V}_{\mathcal{B}}$ of all the balls in $\mathcal{X}$. Therefore, even when the covariate’s dimension is large, the error can be small if the covariate features have strong correlations so that the intrinsic dimension $d$ is small. This is the finding first made by Kpotufe, (2011) on real-valued kNN regression, and we extend it to RKHS-valued kNN regression.
The rate (14) is the same as the minimax optimal rate for estimating a Lipschitz-continuous real-valued regression function when the covariate distribution $P(x)$ has the intrinsic dimension $d$ (Kpotufe,, 2011, Theorem 2). An interesting point is that the same rate is attained with RKHS-valued kNN regression where the output space is an RKHS that can be infinite-dimensional. Similar observations have been made for RKHS-valued kernel ridge regression (Li et al.,, 2022; 2024).
@@
4 Experiments
We describe experiments to asses the empirical performance of kNNSampler in recovering the distribution of missing values. Section 4.1 explains the settings, evaluation metrics and benchmark methods. Section 4.2 describes and discusses the results.

### v2

@@ Abstract
We study a missing-value imputation method, termed kNNSampler, that imputes a given unit’s missing response by randomly sampling from the observed responses of the $k$ most similar units to the given unit in terms of the observed covariates. This method can sample unknown missing values from their distributions, quantify the uncertainties of missing values, and be readily used for multiple imputation. Unlike popular kNNImputer, which estimates the conditional mean of a missing response given an observed covariate, kNNSampler is theoretically shown to estimate the conditional distribution of a missing response given an observed covariate. Experiments illustrate the performance of kNNSampler. The code for kNNSampler is made publicly available. [^1] https://github.com/SAP/knn-sampler
@@ 1 Introduction
Missing values occur in real-world datasets for various reasons, such as non-response in surveys and sensor failures. Imputation — filling in missing values with their estimates — is a common preprocessing step used to address missing data. Over the decades, various imputation methods have been proposed, ranging from simple statistical techniques to machine learning algorithms (e.g., rubin1976inference; schafer1997analysis; schafer2002missing; little2002statistical; mattei2019miwae; enders2022applied).
kNNImputer (troyanskaya2001missing) is one of the most widely used imputation methods, owing to its simplicity and availability in popular software packages such as scikit-learn [^2] https://scikit-learn.org/stable/modules/generated/sklearn.impute.KNNImputer.html (scikit-learn). It imputes a missing response variable (e.g., customer satisfaction level) of a given unit (e.g., a customer) as the average of the observed responses of the $k$ most similar units to the given unit in terms of observed covariates (e.g., age, gender, occupation). This is to predict the missing response by $k$ nearest neighbours (kNN) regression (stone1977consistent) so the imputation is an estimate of the conditional expectation of the missing response given a covariate. The method has been widely used in science and engineering, and many extensions have been proposed (e.g., garcia2009k; tutz2015improved; de2016missing; huang2017cross; faisal2021multiple).
@@
An issue of kNNImputer, shared by other regression-based imputers, is that the distribution of imputations can be significantly different from the distribution of true (hidden) missing values. This is because, as mentioned, an imputation of kNNImputer is an estimate of the conditional expectation of a missing response, thus tending to be a deterministic function of the covariate. As a result, the distribution of imputed responses is concentrated around the regression curve, even when the distribution of missing responses has large variability. This is illustrated in Figures 1 and 2, where the true conditional distribution of a missing response is bimodal when the covariate is small, but the distribution of imputations is unimodal and many imputations take values never realized by the true missing values. A substantial bias can occur in an analysis of such a distorted imputed dataset, for example, when estimating the variance, quantiles and modes in the population. (See Sections 2.1 and 2.2 for more formal discussions.)
@@
kNNSampler can be interpreted as an instance of hot deck, classic imputation methods widely used in practice for socio-economic and public health surveys, including the U.S. Census Bureau’s Current Population Survey and the National Center for Education Statistics (e.g., andridge2010review). In a hot deck method, a missing value of a given unit is imputed as one of the response values of the units belonging to the same “adjustment cell” as the given unit. The method is called random hot deck if the imputation is selected randomly from the adjustment cell; it is called nearest-neighbour hot deck if nearest neighbours define the adjustment cell (little2002statistical, Example 4.9). kNNSampler is thus essentially a nearest-neighbour random hot deck method. However, while classic and widely used, hot deck methods have not been well established theoretically (andridge2010review).
@@
  • We derive an error bound between the kNN and true conditional distributions for any given, fixed covariate, in terms of the number $n$ of observed response-covariate pairs, the number $k$ of the nearest neighbours, and other problem-specific constants. The error is measured by the maximum mean discrepancy (MMD) (gretton2012kernel), a distance metric on probability distributions that metrizes the weak convergence (simon2023metrizing), between the kNN and true conditional distributions. It holds under a Lipschitz condition that the response’s conditional distribution changes smoothly when the covariate changes continuously. A consequence of the bound is the statistical consistency of the kNN conditional distribution, in that the error decreases to zero as the sample size $n$ goes to infinity, if the number $k$ of nearest neighbours increases to infinity at a rate slower than $n$. This offers a theoretical foundation of the kNNSampler and thus the nearest-neighbour random hot deck.
  • To derive the bound, we analyze the mean embedding of the kNN conditional distribution in a reproducing kernel Hilbert space (RKHS) as a novel estimator of the mean embedding of the true conditional distribution, known as conditional mean embedding (muandet2017kernel, Chapter 4), which is the RKHS-valued regression function (grunewalder2012conditional). The RKHS distance between these two embeddings is equivalent to the MMD between the kNN and true conditional distributions. Our bound leads to the consistency and convergence rates for the novel kNN-based estimator of the conditional mean embedding.
  • Our analysis extends the error analysis by kpotufe2011k on real-valued kNN regression to RKHS-valued regression in which the response variable is infinite-dimensional. As a byproduct, we prove that the required sample size to attain a given level of precision increases exponentially not with the covariate’s ambient dimension but with the intrinsic dimension of the covariate distribution. Therefore, the kNNSampler may not be severely affected by the curse of dimensionality if the covariate distribution has a low intrinsic dimension.
This paper is organised as follows. We describe the proposed approach in Section 2 and its theory in Section 3. We report experimental results on synthetic data in Section 4 and on real solar-power data in Section 5.
@@ 2.1 Setting
This assumption implies that the probability of a unit missing its response is determined by the unit’s covariate and is not affected by the response. Therefore, it is an instance of the Missing-At-Random (MAR) assumption (rubin1976inference). In the special case where the covariate distributions for the two cases are the same, Q(x~)=P(x~)Q($\tilde${x})=P($\tilde${x}), the assumption can be interpreted as the Missing-Completely-At-Random (MCAR) assumption where missingness occurs completely randomly.
Under this setup, missing responses may be imputed by estimating the unknown conditional distribution $P(y|x)$ of a response given a covariate and sampling from it. This is what the kNNSampler does.
Remark 1.
Sampling imputations from the conditional distribution is needed for unbiased estimation of a quantity of interest and its well-calibrated uncertainty quantification. We informally describe this from the Bayesian perspective of rubin1987multiple, where the quantity of interest, denoted here by $\theta$, is treated as a random variable. For the frequentist perspective, see rubin1987multiple; Rubin96; murray2018multiple. A Bayesian analysis is done by computing the posterior distribution of $\theta$ given the observed data in (1):
$P(\theta\mid\mathcal{D}_{n},~\tilde{x}_{1},\dots,\tilde{x}_{m})$
$=\int\underbrace{P(\theta\mid\mathcal{D}_{n},~(\tilde{x}_{1},\tilde{y}_{1})\dots,(\tilde{x}_{m},\tilde{y}_{m}))}_{(A)}\cdot\underbrace{P(\tilde{y}_{1},\dots,\tilde{y}_{m}\mid\mathcal{D}_{n},~\tilde{x}_{1},\dots,\tilde{x}_{m})}_{(B)}d\tilde{y}_{1}\cdots d\tilde{y}_{m},$
where $(A)$ is the posterior distribution of $\theta$ given observed dataset $\mathcal{D}_{n}=\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ and imputed dataset $(\tilde{x}_{1},\tilde{y}_{1}),\dots,(\tilde{x}_{m},\tilde{y}_{m})$ and is computed by a standard Bayesian analysis, treating the imputations as observed data; $(B)$ is the conditional distribution of missing responses $\tilde{y}_{1},\dots,\tilde{y}_{m}$ given $\mathcal{D}_{n}$ and observed covariates $\tilde{x}_{1},\dots,\tilde{x}_{m}$ for missing responses, and can be written as
$P(\tilde{y}_{1},\dots,\tilde{y}_{m}\mid\mathcal{D}_{n},~\tilde{x}_{1},\dots,\tilde{x}_{m})=\prod_{i=1}^{m}P(\tilde{y}_{i}|\tilde{x}_{i}),$
[…]
@@ 2.2 Issue with kNNImputer and Regression-based Imputers
Before describing the proposed kNNSampler, we discuss an issue with the widely used kNNImputer (troyanskaya2001missing) and other regression-based imputation methods.
@@
kNNImputer (troyanskaya2001missing) imputes the missing response $\tilde{y}_{i,{\rm miss}}$ of the unit with observed covariate $\tilde{x}_{i}$ as the average of the observed responses $y_{j_{1}},\dots,y_{j_{k}}$ of its $k$-nearest neighbors $x_{j_{1}},\dots,x_{j_{k}}$:
@@
This is kNN regression (e.g., gyorfi2002distribution) and thus estimates the conditional mean of the missing response $\tilde{y}_{i,{\rm miss}}$ given the observed covariate $\tilde{x}_{i}$:
@@
Remark 2.
Consider the Bayesian analysis in Remark 1. If the missing responses are imputed by deterministic, regression estimates $\hat{y}_{1,{\rm imp}},\dots,\hat{y}_{m,{\rm imp}}$, the posterior distribution becomes that of the quantity of interest $\theta$ given observed and imputed datasets, both treated as observed:
$P(\theta\mid\mathcal{D}_{n},~(\tilde{x}_{1},\hat{y}_{1,{\rm imp}})\dots,(\tilde{x}_{m},\hat{y}_{m,{\rm imp}}))$
This ignores uncertainties in the missing responses, leading to final uncertainty estimates for $\theta$ that are overconfident (a prediction interval may be much narrower than the actual error of a point estimate).
@@
The number of nearest neighbors $k$ is a hyperparameter of kNNSampler. The theoretical and empirical results below indicate that $k$ should not be fixed to a prespecified value (e.g., $k=5$), and should be chosen depending on the available data. One way is to perform cross-validation for kNN regression on the data $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ and select $k$ among candidate values that minimizes the mean-square error on held-out observed responses, averaged over different training-validation splits. In particular, the present work uses Leave-One-Out Cross-Validation (LOOCV) using the fast computation method recently proposed by kanagawa2024fast.
@@ 2.5 Multiple Imputation with kNNSampler
An analysis can then be made based on the standard procedure of multiple imputation (rubin1987multiple).
@@ 3 Theory
We describe a theory for kNNSampler’s conditional distribution (6) as an estimator of the true conditional distribution. We shall show that, as the number $k$ of nearest neighbors increases at an approximate rate as the increase of the number $n$ of observed covariate-response pairs, the kNN conditional distribution converges to the true one in the Maximum Mean Discrepancy (MMD) (gretton2012kernel), which implies the convergence in distribution (sriperumbudur2010hilbert, Section 5). We prove this by adapting the proof of the convergence rates of real-valued kNN regression by kpotufe2011k to Hilbert space-valued kNN regression. [^4] Hilbert space-valued kNN regression was also analyzed in lian2011convergence, but their results are not directly applicable to our case. This is because lian2011convergence assumes that Hilbert space-valued noises are independent of input variables, but this assumption is too strong in our case.
We use the framework of kernel mean embedding (muandet2017kernel) in which every probability distribution is represented as a distinct point in an infinite-dimensional feature space known as a reproducing kernel Hilbert space (RKHS). The true and kNN conditional distributions are represented as points in an RKHS, and the distance between them, which is the MMD, quantifies the estimation error. An upper bound on this distance is obtained in terms of the sample size, the number of nearest neighbours, and other relevant quantities.
@@ 3.1 RKHS Embeddings of Conditional Distributions
This is called reproducing kernel of the RKHS. The RKHS and the reproducing kernel are one-to-one, so an RKHS can be induced by defining a kernel. For example, if $\mathcal{Y}=\mathbb{R}^{p}$, the Gaussian kernel $\ell(y,y^{\prime})=\exp(-\alpha\|y-y^{\prime}\|^{2})$ for $\alpha>0$ is the reproducing kernel of a certain RKHS $\mathcal{H}$, and there exists an infinite-dimensional feature map $\Phi$ that induces the Gaussian kernel as (7). See e.g. SteChr2008; kanagawa2025gaussian for details on RKHSs.
@@
In this case, the RKHS is called characteristic (sriperumbudur2010hilbert). For example, Gaussian, Matérn and Laplace kernels induce characteristics RKHSs.
@@
The mean embedding of the conditional distribution is known as conditional mean embedding (song2009hilbert; song2013kernel) and its estimator based on a regularized least-squares algorithm has been studied extensively (e.g., grunewalder2012conditional; li2022optimal; li2024towards). The mean embedding of the kNN conditional distribution in (8) is a new estimator of the conditional mean embedding. Its analysis below is thus a new contribution to the RKHS literature and may be of independent interest.
@@ 3.2 Assumptions
We describe key assumptions for the analysis, which follow kpotufe2011k with appropriate modifications.
The conditional mean embedding in (8) is the conditional expectation of the response feature vector $\Phi(y)$ given a covariate $x\in\mathcal{X}$; thus, it is the RKHS-valued regression function (grunewalder2012conditional). We assume that the map from a covariate $x$ to the conditional mean embedding $\Phi(P(\cdot\mid x))$ is smooth in the sense that it is Lipschitz continuous.
@@
This assumption is satisfied, for example, if $\mathcal{X}=\mathbb{R}^{p}$ with $p\geq 1$, in which case $\mathcal{V}_{\mathcal{B}}\leq p+2$ (e.g., mohri2018foundations, Exercise 3.17).
@@ 3.3 Error Bounds and Convergence Rates
Under the above assumptions, the distance between the true and kNN conditional distributions can be upper-bounded as follows. The proof, provided in Appendix A, is an adaptation of the proof of kpotufe2011k, which is an upper error bound on real-valued kNN regression, to our setting of RKHS-valued regression.
@@
$\left\|\Phi(P(\cdot\mid x))-\Phi(\hat{P}(\cdot\mid x))\right\|_{\mathcal{H}}^{2}\leq C_{1}\frac{\ln(n)}{k}+C_{2}\left(\frac{k}{n}\right)^{2/d},$ (12)
@@
On the other hand, if $k$ is fixed to a constant value (e.g., $k=1$), the variance term does not decrease even if the sample size increases. These observations are well known for real-valued kNN regression (e.g., gyorfi2002distribution).
@@
The above consistency (13) implies the convergence in distribution (or weak convergence) of the kNN conditional distribution $\hat{P}(\cdot\mid x)$ to the true one $P(\cdot\mid x)$ if the response space $\mathcal{Y}$ is a compact metric space (e.g., $\mathcal{Y}$ is a bounded closed subset of an Euclidean space) and $\mathcal{H}$ is a universal RKHS [^6] An RKHS $\mathcal{H}$ consisting of functions on a metric set $\mathcal{Y}$ is called universal if any continuous bounded function $f:\mathcal{Y}\to\mathbb{R}$ can be approximated arbitrarily well in terms of the supremum norm by functions in $\mathcal{H}$., such as the RKHSs of Gaussian, Matérn and Laplace kernels (sriperumbudur2010hilbert, Theorem 23); see simon2023metrizing for more generic conditions. That is, under these conditions, the expectation of any continuous bounded function $f:\mathcal{Y}\to\mathbb{R}$ under the kNN distribution $\hat{P}(\cdot\mid x)$ converges to the expectation under the true distribution $P(\cdot\mid x)$:
@@
The rate (14) shows that the required sample size $n$ to attain a desired error level increases exponentially with respect to the intrinsic dimension $d$ of the covariate distribution $P(x)$, not the ambient dimension of the input space $\mathcal{X}$, which is captured by the VC dimension $\mathcal{V}_{\mathcal{B}}$ of all the balls in $\mathcal{X}$. Therefore, even when the covariate’s dimension is large, the error can be small if the covariate features have strong correlations so that the intrinsic dimension $d$ is small. This is the finding first made by kpotufe2011k on real-valued kNN regression, and we extend it to RKHS-valued kNN regression.
The rate (14) is the same as the minimax optimal rate for estimating a Lipschitz-continuous real-valued regression function when the covariate distribution $P(x)$ has the intrinsic dimension $d$ (kpotufe2011k, Theorem 2). An interesting point is that the same rate is attained with RKHS-valued kNN regression where the output space is an RKHS that can be infinite-dimensional. Similar observations have been made for RKHS-valued kernel ridge regression (li2022optimal; li2024towards).
@@
4 Synthetic Data Experiments
We describe experiments to assess the empirical performance of kNNSampler in recovering the distribution of missing values. Section 4.1 explains the settings, evaluation metrics, and benchmark methods. Section 4.2 describes and discusses the results.